OpenStudy (anonymous):

Factory completely. a) 3x2 -7x - 6 b) 4x2 - 9y2

OpenStudy (anonymous):

For a use the quadratic formula.

OpenStudy (anonymous):

For question (a) i mean.

OpenStudy (solomonzelman):

\[a)~~~(3x+2)(x-3)\]

OpenStudy (solomonzelman):

\[b)~~~~4x^2-9y^2=(2x)^2-(3y)^2=(2x+3y)(2x-3y)\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

can you show me the next step with a?

OpenStudy (solomonzelman):

The a and b I posted, are factored completely.

OpenStudy (anonymous):

so as the answer what would i writ?

OpenStudy (solomonzelman):

\[a)~~~~~~(3x+2)(x-3)~~~~~~~~and~~~~~~~~b)~~~~~~(2x+3y)(2x-3y)\]

OpenStudy (anonymous):

can you please write down how you did it step by step?

OpenStudy (anonymous):

@SolomonZelman please

OpenStudy (solomonzelman):

I can't explain the a, I used my drain. I am not kidding my thing process was. \[(3x-+~~~~~~)(x-+~~~~~~)~~~~~~~~~~~~plunging~~~i n~~factors~~of~~-6.\]

OpenStudy (solomonzelman):

(not drain, brain)

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

Do you need more work on the b ?

OpenStudy (anonymous):

that would be great.

OpenStudy (solomonzelman):

\[4x^2-9y^2\]WE KNOW,\[4x^2=(2x)^2~~~~~~~~~~and~~~~~~~~~~9y^2=(3y)^2\]SO, RE-WRITE,\[(2x)^2-(3y)^2\]then I used\[a^2-b^2=(a+b)(a-b)\]